[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Oct 21, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Oct 22 03:16:24 UTC 2017


HSU Shift Log for Sat Oct 21 2017 23:16:19 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: HSU phones operational as of 2200 EDT; SDM called  
to confirm.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171022/4e64d7c5/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list